OpenStudy (anonymous):

What is the average of positive integers from 1 to 100 inclusive? (A) 49 (B) 49.5 (C) 50 (D) 50.5 (E) 51

OpenStudy (anonymous):

Aslam has 4 times as many books as Salman and 5 times as many as Javed. If Javed has more than 40 books, what is the least number of books that Aslam could have? (A) 200 (B) 205 (C) 210 (D) 220 (E) 24

OpenStudy (welshfella):

average = (1+100) / 2

OpenStudy (welshfella):

Javed has more than 40 books , so least he can have is 41 so aslam would then have 41* 5 = 205 books but aslam has 4 times as many books as Salman so 205 cannot be the correct amount So what do you think we should try next?

OpenStudy (welshfella):

the number of books that aslam has must be divisible by 4 and 5

OpenStudy (anonymous):

Thank You! Got it.

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

220 is the right answers of books that Aslam could have but I could not find the exact 220

OpenStudy (welshfella):

yea 220 is correct

OpenStudy (anonymous):

Other than 205 there are 2 options,So y selecting 220

OpenStudy (anonymous):

Can U plz tell me that how 220 is correct

OpenStudy (welshfella):

javed would have 44 44 8 5 = 220 for aslam 220 / 4 = 55 Salman would have 55 books

OpenStudy (welshfella):

* 44 * 5

OpenStudy (welshfella):

220 is the only number that is divisible by 4 and 5 (except for 200 but it cant be 200 because javed must have more than 40 books)

OpenStudy (welshfella):

- and , of course, the number of books must be a whole number.

OpenStudy (welshfella):

is that clear? I can help further if you want.

OpenStudy (anonymous):

Yes clear now,thanks

OpenStudy (welshfella):

yw